OpenStudy (anonymous):

Help! How would I solve this??f(x)=2x+6/x+2, and f(6)

OpenStudy (anonymous):

2(6)+6=18 6+2= 8 18/8 = 9/4

OpenStudy (anonymous):

@lex245 could you help me with one more? f(x)=2x+6/x+2

OpenStudy (anonymous):

yeah what is it

OpenStudy (anonymous):

F(x)=2x+6/x+2 if f(x)=1

OpenStudy (anonymous):

2(1) +6 = 8 1+2= 3 so it would equal 8/3

OpenStudy (anonymous):

Awesome! Thank you so much.

OpenStudy (anonymous):

you would just plug in the number they give you for x and then just divide regularly

OpenStudy (anonymous):

cool, glad i could help!